Jiawei's age is twice that of his brother's. His father is 5 times of his. If their total age is at most 55 years,what is his father's maximum possible age?
1
Expert's answer
2020-05-11T17:57:14-0400

"x\;-\;age\;of\;brother\\\;\;2x\;-\;age\;of\;Jiawei\;\;\\10x\;-\;age\;of\;father\;\\\;10x\;+\;2x\;+\;x\;=\;13x\leq\;55\;\\\;x\;\leq\;4.2307\;\;\\10x\;\leq\;42.307\;\\\;So,\;maximum\;possible\;age\;of\;his\;father\;is\;42\;\\because\;it\;should\;be\;a\;whole\;number."